### v1

# The Value of AI Advice: Personalized and Value-Maximizing AI Advisors Are Necessary to Reliably Benefit Experts and Organizations

Abstract
Despite advances in AI’s performance and interpretability, AI advisors can undermine experts’ decisions and increase the time and effort experts must invest to make decisions. Consequently, AI systems deployed in high-stakes settings often fail to consistently add value across contexts and can even diminish the value that experts alone provide. Beyond harm in specific domains, such outcomes impede progress in research and practice, underscoring the need to understand when and why different AI advisors add or diminish value. To bridge this gap, we stress the importance of assessing the value AI advice brings to real-world contexts when designing and evaluating AI advisors. Building on this perspective, we characterize key pillars – pathways through which AI advice impacts value – and develop a framework that incorporates these pillars to create reliable, personalized, and value-adding advisors. Our results highlight the need for system-level, value-driven development of AI advisors that advise selectively, are tailored to experts’ unique behaviors, and are optimized for context-specific trade-offs between decision improvements and advising costs. They also reveal how the lack of inclusion of these pillars in the design of AI advising systems may be contributing to the failures observed in practical applications.
Keywords AI-advised decision-making  $\cdot$ human-AI collaboration  $\cdot$ rule sets  $\cdot$ high-stakes decisions
1 Introduction
Artificial intelligence (AI) systems are poised to play an increasingly vital role in augmenting human decision-makers across high-stakes environments, such as in healthcare [1]. Yet, despite advances in AI’s generalization performance and interpretability, AI systems deployed to advise decision-makers in critical settings fail to consistently improve outcomes across contexts and expert behaviors [2, 3, 4, 5, 6, 7, 8]. For example, AI models with generalization accuracy that is comparable to or even exceeds that of human experts are increasingly used to advise radiologists in diagnostic tasks [5, 8, 3]. Yet, these tools have been shown to fall short of expectations, sometimes leading experts to make poorer decisions or prolonging the decision-making process [5, 9, 3] without adding sufficient value – outcomes which have also led to the under-utilization or abandonment of the tools after deployment [5].
The lack of benefit provided by AI tools is widespread across domains, task types, and human behaviors. A recent survey found that 56 papers on AI-assisted decision-making published between January 2020 and June 2023 reported that AI-assisted humans produced worse outcomes relative to the human alone [10]. Such detrimental outcomes run the risk of reducing trust in AI advice and impeding progress towards AI’s positive impact and adoption in critical decision-making domains. Nevertheless, to date, research on the development [10, 11, 12, 13] and evaluation [3, 14] of AI in AI-advised decision-making settings is often focused on improving AI’s state-of-the-art performance according to measures such as the AI’s accuracy on decision tasks without appropriate regard for how the AI’s advice will heterogeneously add to or diminish the value of the individuals it is meant to advise. To our knowledge, no existing work offers a comprehensive framework necessary to understand and mitigate the failures of AI advising systems seen in practice. Such a framework is critical to anticipating and explaining when AI advice will enhance or diminish the value provided by the individuals it aims to support. It is also necessary to guide the design of reliable, value-adding AI advisors. This paper seeks to address this gap.
We focus on a common high-stakes setting in which a human expert first makes an initial judgment on the best decision to take, and an AI may subsequently offer advice to the expert on the best decision to take before the expert finalizes their decision [5]. Henceforth we refer to this setting as AI-assisted Decision-Making (AIaDM), and refer to the AI as the AI advisor.
We propose that enhancing the positive impact of AI advisors in practice requires a fundamental shift toward a system-level, value-based approach to guide their design, learning, and evaluation. Consequently, the focus of this work is not the design of a new AI method with superior decision-making performance. First, we propose that AI advisors’ goal is to reliably offer the most added-value to their context, and we outline key pathways — referred to as pillars — through which AI advice can impact value in AIaDM contexts and are thereby integral to achieving the AI advisor’s goal. Building on the objective and pillars, we develop a framework that can be used to produce Reliable and Value-maximizing AI (ReV-AI) advisors, and we use our framework to develop a concrete proof-of-concept ReV-AI method – referred to as TeamRules (tr) – for generating inherently interpretable rule-based advisors, each personalized to an individual decision-maker and context. Using real world data along with simulation grounded in human behavior, we asses the value that different AI advisors’ provide across contexts, explore how different pillars affect AI advisors’ value impact, and shed light on the distinct advising strategies that ReV-AI advisors utilize to reliably add value.
We find that ReV-AI advisors produce consistent benefits across contexts and human behaviors by producing distinctly advantageous advising behaviors tailored to their context and advisee. Crucially, we also find that accounting for all of the pillars is key to yielding consistent benefits. Finally, our results reveal the risks posed by AI advisors which do not account for some of the pillars or elements of the framework, not only demonstrating how AI advisors can diminish value in some contexts, as has been shown to occur in practice, but also offering actionable insight into why such outcomes arise.
The AI Advising Objectives: Offer Reliable, Value-Maximizing Advice
Perhaps the most fundamental objectives of an AI advisor are that (a) it learns to produce advice that maximizes the value it adds to its context relative to what can be achieved by the human alone and that (b) an AI advisor is reliable, such that when there is no opportunity to provide value-adding advice, it does not diminish the value delivered by the expert alone. Crucially for our work, the value added by an AI advisor encompasses not only the improvement in the decisions made by the human-AI team but also the practical costs associated with advising the human. Henceforth, we refer to advisors aiming to achieve these objectives as Reliable and Value-Maximizing AI (Rev-AI) advisors.
Pillars for Developing AI Advisors
Central to our proposed system-level, value-based approach are pillars that characterize key pathways through which AI advice can add or diminish value in AIaDM settings and are thereby foundational for producing Rev-AI advisors and to our understandings of failures across applications in practice.
  • The overlooked costs of AI advice. Recent studies on AI-advised classification tasks, such as medical diagnosis, often focus on optimizing the AI’s standalone predictive performance [11, 12, 13]. However, for AI advisors to add value in practice, improving predictive performance alone is insufficient. For instance, while double-reading diagnostic images by radiologists improves accuracy, it has not been widely adopted in the U.S. due to cost-effectiveness concerns [15, 16, 17]. Thus, to develop AI advisors that add value to their context, beyond improving decision outcomes, it is essential to account for the costs associated with integrating AI advice into the decision-making process.
    A significant yet overlooked cost in the development and evaluation of AI advisors is the burden placed on experts when they engage with AI advice. AI advice can only improve an expert’s independent judgment when it contradicts it. Further, evidence shows that decision-makers in high-stakes settings experience significant time costs when reconciling AI advice that contradicts their judgment [18, 19, 9, 5]. Similar to previous innovations deemed cost-ineffective, excessive human engagement costs from reconciling disagreements with AI advice have been shown to impede the use of AI advice altogether [5, 20]. Advancing AI advisors’ positive impact on practice thus hinges on developing frameworks that produce cost-effective AI advisors, accounting for the human costs of reconciling contradictory advice. While crucial to the value proposition of AI advisors in critical settings, the human engagement and its implications have not been considered thus far in practice or in research on AI advising methods. Importantly, the value added by AI advice depends on these costs, and the cost-effectiveness of advice is moderated by the idiosyncratic trade-off within each context between added advising costs and the value generated from changes in the final decisions—which we henceforth refer to as the context’s trade-off.
  • Humans are idiosyncratic and imperfect advice-takers.
    Human discretion in accepting or rejecting contradictory AI advice is both idiosyncratic and imperfect: individuals may reject accurate advice or accept incorrect recommendations [21, 22, 23, 24, 25, 26, 7, 8, 14]. We refer to this tendency as an individual’s Algorithm Discretion Behavior (ADB). ADB affects the cost-effectiveness of AI advice: given the human’s engagement to reconcile contradictory AI advice is costly, a higher likelihood of rejecting AI advice reduces its cost-effectiveness.
  • Whether advice is accepted depends on the advice itself: offer inherently convincing advice.
    An individual’s likelihood of accepting AI advice varies across decision instances and depends on properties inherent to the advice, such as its confidence [27, 28, 29]. Therefore, complementing AI recommendations with inherent properties of the advice has become crucial in high-risk settings, as such properties truthfully characterize the AI’s recommendation [30, 31]. We refer to inherent properties of AI advice that impact the likelihood with which an individual will accept it as ADB-bound information. Crucial for learning cost-effective advice, research shows that the human’s likelihood of accepting AI advice is influenced by the AI’s reported confidence in addition to the human’s self-reported (and potentially miscalibrated) confidence in their own decision [27, 28, 29]. Hence, developing value-maximizing AI advisors hinges on learning and offering advice that is simultaneously sufficiently inherently persuasive and accurate to add value, per the context’s trade-off.
  • Humans are idiosyncratic and imperfect decision-makers. Lastly, the expert’s decision-making behavior is central to the development of cost-effective AI advisors. Due to the idiosyncratic strengths and weaknesses of experts’ judgment, the value of AI advice also depends on whether the AI’s recommendation is more likely to be correct than the human’s independent decision [32].
Framework for Producing Reliable Value-Maximizing AI Advisors
Figure 1: a) Framework for producing reliable, value-maximizing AI advisors. b) TeamRules + Human Partner Decision Making Process
We build on the pillars to develop a general framework for producing ReV-AI advisors. In a departure from existing AI advisors, the ReV-AI framework constitutes a system-level and value-based approach for developing and evaluating AI advisors. ReV-AI advisors selectively offer advice that maximizes added value while avoiding actions that could diminish the value they provide to the expert within their context. Crucially, the advisors are grounded in the key pillars through which AI advice can impact the context’s value. Below we briefly note the key elements of the ReV-AI Learning Framework visualized in Figure 1a.
Nature of advice: decision recommendation and inherent ADB-bound information. A ReV-AI advisor provides both a decision recommendation and ADB-bound information that influences the likelihood of its advice being accepted by the human advisee. We consider advisors that provide confidence—the estimated likelihood that the advice is correct—as the complementary ADB-bound information offered. However, our framework can also generate recommendations alongside other inherent properties that inform the human’s ADB and, by extension, the value of the advice. [^1] We focus on confidence here because it allows us to simulate diverse behaviors by building on prior work which has established with real human interactions how AI’s confidence along with the human’s self-assessed confidence affect the human’s receptiveness of advice [27].
Learning to offer value-maximizing advice personalized to a given human and context. Our framework includes end-to-end learning of a model, which constitutes a repertoire of decision recommendations with corresponding confidence levels that are offered selectively during deployment, so as to add the most value across the decision space. Note that to add the most value to a context, a ReV-AI advisor can learn to produce advice that adds significant value for some decisions at the possible expense of accuracy or persuasiveness on other decisions. Value-maximizing advice is learned through an objective grounded in the pillars, which reflect how AI counsel either enhances or diminishes value in its context. The value of the advice is thus influenced by (1) its inherent persuasiveness given the human’s confidence and corresponding ADB, (2) its likelihood of being correct relative to the likelihood the human’s independent decision is correct, and (3) the context’s trade-off between the human’s engagement with the AI recommendation and the decision benefits of the advice. Hence, producing value-maximizing advice grounded in the ReV-AI advising objective requires data on the human’s expertise, confidence, and ADB. We discuss data collection and related modeling in the Methods section.
Selective advising. In some cases, even the best advice learned by a ReV-AI advisor may not necessarily be cost-effective enough to add value. Thus, a crucial component of a ReV-AI advisor is a selective advising mechanism that allows it to withhold advice when the best learned advice is expected to diminish value. This mechanism is integrated into the learning process, ensuring that the advice repository is optimized to add value while accounting for the selective withholding of advice if the advice is not expected to contribute positively.
TeamRules: An Inherently Interpretable ReV-AI Advisor.
The ReV-AI learning framework can be used to derive different concrete methods for generating ReV-AI advisors of different model classes. In this work, we develop and evaluate a rule-based prototype algorithm, TeamRules (tr), which produces ReV-AI advisors that generate decision recommendations in the form of inherently interpretable rules. The tr advising process during deployment is shown in Figure 1b. Inherently interpretable models are especially suitable for and are commonly used to advise humans in high-stakes contexts [31, 33, 34, 35, 36, 37, 38, 39] because they faithfully represent the model’s actual decision-making process [40].
2 Results
Table 1: Summary of advisor generating methods, human behavior simulations, and measures of performance
Accounts for:
Advising
Framework
Advises
Selectively?
Human
Expertise?
Human
ADB?
Cost-benefit
Trade-off?
TR
TR-no(ADB)
TR-no(Cost)
TR-no(ADB, Cost)
Task-Only (Current Practice)
(a) Advisor Generators
Simulated Expertise Behaviors Simulated Confidence Behaviors Difficulty-Biased Decisions Group-Biased Decisions Accuracy-Biased ADB Group-Biased ADB Likelihood of human error varies with task difficulty [41]. Likelihood of human error varies across groups in the population [42, 43]. Human’s confidence in their independent judgment varies with but does not perfectly align with their decision accuracy [44, 45, 28]. Human confidence in their independent judgment varies across groups in the population [46].
(b) Types of Simulated Human Expertise and Confidence Behaviors
Metric Definition Value Added (Accuracy Improvement w.r.t. Human) -- (Cost-Benefit Trade-Off)*(Advising Costs Incurred) Accuracy Improvement w.r.t. Human (Number of Human Decisions Corrected Through Advising In Group) $\div$ (Total Number of Decision Instances) Advising Costs Incurred (Number of Pieces of Advice In Group) $\div$ (Total Number of Decision Instances) Advising Costs Incurred (In Accuracy Units [AU]) (Cost-Benefit Trade-Off)*(Advising Costs Incurred) Advising Confidence  [^2] Note that the confidence of rule-based advice corresponds to the precision of the rule. Thus, while the confidence is a measure of the quality of the rules used for advising and estimated prior to deployment, advising accuracy is a measure of the quality of decision advice offered under selective advising. tr selectively offers a rule as advice only if the advice is expected to add value. Hence, rules with any given precision are invoked selectively based on their expected added value, and may yield higher or lower advising accuracy. (Mean Confidence of Contradicting Advice for Instances Advised On In Group) Advising Accuracy (Number of Correct Pieces of Advice In Group) $\div$ (Total Number of Pieces of Advice In Group) Advising Rate (Number of Pieces of Advice In Group) $\div$ (Total Number of Decision Instances In Group) Advice Acceptance Rate (Number of Pieces of Advice Accepted By DM In Group) $\div$ (Number of Pieces of Advice Offered In Group)
(c) Measures of Performance
We conduct experiments to answer a set of key questions. First, we examine the value added by tr and alternative advisors to assess whether any of the them can be relied on to consistently add value across contexts. We also study what, if any, are the risks of ignoring the pillars, and what distinct advising strategies do tr’s advisors produce to reliably add value across contexts. Answering these questions requires consideration of a diverse set of contexts with different kinds of AI advisors produced to advise humans with varied human expertise and discretion behaviors in contexts with different advising cost-benefit trade-offs.
Alternative Advising Frameworks. We evaluated the value added by tr and other rule-based advisors that do not account for subsets of key pillars, including evaluation of the task-only advisor, which follows the current common practice in which AI models are trained to excel at a task by maximizing generalization accuracy and advising the human on all decisions. All advisor-generating frameworks considered, along with which pillars they account for, are shown in Table 1(c). The evaluations aim to reveal the potential performance benefits and risks of alternatives that do not consider all or some of the key pillars.
Figure 2: Value added by different AI advisors relative to independent human for environments with varying trade-offs between advising costs and decision benefits for Heart Disease, FICO, and HR datasets. Value added is the improvement in human’s decisions from advising minus costs incurred from human engagements. X-axis reflects contexts in which the cost from 1/x human engagements are equivalent to the loss from one incorrect decision: higher x-values reflect that fewer human engagements are permissible to achieve a given decision benefit. task-only model generalization accuracy is 0.78 (Heart Disease), 0.71 (FICO), 0.79 (HR). The human’s standalone generalization accuracy for each dataset and decision behavior is 0.71 (Heart Disease Difficulty-Biased), 0.83 (Heart Disease Group-Biased), 0.73 (FICO Difficulty-Biased), 0.70 (FICO Group-Biased), 0.70 (HR Difficulty-Biased), 0.71 (HR Group-Biased). Results show average value-added +- SE (shaded region) over 10 repetitions.
Data. We use three data domains which in practice represent consequential decisions: the Heart Disease Dataset in which labels reflect the presence of heart disease for a patient [47], the FICO dataset [48] reflecting loan applications to inform loan decisions, and an HR employee attrition dataset in which labels reflect employee departure, and thus such data can be used in practice to inform employment retention decisions. [^3] https://www.kaggle.com/datasets/pavansubhasht/ibm-hr-analytics-attrition-dataset, note this is a fictional dataset that represents a real-world task
Simulations. To explore AI advising performance for a wide range of settings, and similarly to recent work that developed methods for human-AI collaboration [49, 32, 50, 51, 52, 53], we simulate diverse human expertise [54, 49, 55, 53], confidence, and corresponding ADB [44, 45, 28, 46] informed by prior work on human behavior, summarized in Table 1(c). Complete details of the simulation procedures and parameters are outlined in Appendix A. We build on prior work that modeled and validated human ADB with real human interactions [27], and which maps the human’s self-reported confidence and the AI advisor’s confidence to the human’s likelihood to accept the AI’s advice. In general, humans are more likely to accept AI advice the higher the AI’s confidence is relative to their own confidence in their decision. We simulate various human confidence patterns from previous studies [44, 45, 28], which lead to different probabilities of accepting AI advice based on the AI’s confidence level for each instance.
Evaluations. For different contexts characterized by different data domains, human expertise and ADB, Figure 2 shows the AI advisors’ added-value relative to the human’s as a function of the context’s trade-off: the cost incurred from reconciling one disagreement is equivalent to the loss from $x$ decision errors. Thus, higher values on the x-axis reflect contexts in which fewer human engagements are tolerated to achieve a given decision benefit, where a human engagement refers to the human reconciling AI advice for a single instance, and AI advice refers to an AI decision recommendation that contradicts the human’s initial judgment. The dashed horizontal line is the value from the human’s independent decisions. All evaluation metrics are summarized in Table 1(c) and discussed in detail in Appendix A.4.
The value-added and lost by AI advisors.
As demonstrated across diverse contexts in Figure 2, tr’s advisors consistently add the most value and, where not superior, perform comparably to alternatives. [^4] In Appendix C.1 and C.2, we provide results for additional settings with even more inhibitive advising cost-benefit trade-offs and with different decision loss cost structures. tr advisors can also be relied on to not diminish the value offered by the human alone. Importantly, we see that tr’s performance is consistent even in contexts where all other AI advisors struggle to add value. For example, it is more difficult to add value in contexts where fewer human engagements are tolerated to achieve a given improvement in decision outcomes (higher x-values), reflecting contexts such as emergency medicine. Yet, in such contexts, tr advisors add the most value and avoid losses from cost-ineffective human engagement. It is also challenging for AI advisors to add value in typical expert settings where the expert’s strong decision-making abilities and confidence in their own judgment limit opportunities for improvement, given this entails producing even more accurate and highly persuasive advice. For example, in Figure 2b, the task-only AI advisor always diminishes value, even when the context incurs no costs from human engagement (x-axis value of 00), demonstrating the model class’s difficulty of learning advice that is superior and convincing enough to improve the human’s decisions at all, let alone cost-effectively when advising costs are less tolerated. To benefit such contexts it is especially crucial to learn and selectively offer value-maximizing advice that is simultaneously sufficiently accurate and convincing to warrant the human’s attention, as demonstrated by tr.
Figure 3: Case Study - Outcomes: Advising outcomes when tr, task-only, tr-no(adb) advisors advise an over-confident expert. The context’s trade-off is $0.1$, reflecting that reconciling up to 10 contradictory pieces of AI advice is deemed cost-effective if the contradictions yield at least one improved decision. Advising costs incurred shown are converted to units of decision loss given the trade-off (AU). The expert has an independent decision accuracy of 90% on the Male population (majority), a 60% accuracy on the Female population (minority), and an overall accuracy of 87.5% on the entire population. This expert is highly confident (97.5% confidence) on the entire population. Results are averages over 20 repetitions $\pm$ SE (vertical line at center of each bar).
Figure 4: Case Study - Advising Strategies: Properties of advice corresponding to Figure 3. Advising overconfident expert leads TR to learn and selectively offer higher accuracy and more confident advice that the expert is more likely to accept, leading to more improvements in decisions from fewer expert engagements. Results from 20 repetitions.
Our results also reveal that AI advisors that do not advise selectively or are not informed by one or more of the pillars not only fail to maximize their value but also risk diminishing the value of the context in which they operate. Task-only advisors, which are commonly deployed in practice, incur the greatest losses, even when their generalization accuracy exceeds the human’s independent accuracy, highlighting the risks associated with the most prevalent AI advisors, and consistent with evidence from fields like radiology, where increased human engagement to reconcile disagreements has often been deemed cost-ineffective [5]. Other advisors, such as tr-no(adb) which includes all but one of tr’s pillars and consequently does not attempt to learn advice that is sufficiently inherently convincing to the human, can also diminish the value offered by the human alone when the context is not tolerant of unlimited human engagements (high x-axis values in Figures 2b,d,e,f,g,h).
We find that when neglecting the human’s ADB, tr-no(adb) matches tr’s value when unlimited human engagements are allowed to achieve a given decision improvement (x-axis value of 0). However, as the context permits fewer engagements to achieve a given decision benefit (higher x-axis values), accounting for the human’s ADB becomes essential to add value. This is because, to be cost-effective, advice must not only be superior but also sufficiently convincing to justify the cost of engagement. Accounting for the human’s ADB is essential to accurately estimate the likelihood of the advice being convincing and accepted. We also find that accounting for the human’s decision-making behavior can be beneficial even in contexts allowing unlimited human engagement. This is evidenced by the relative value added by tr-no(adb,cost) (which considers the human’s expertise) compared to task-only. The advantage of tr-no(adb,cost) stems from its ability to focus on decisions the human is less likely to handle correctly and to withhold advice when it is less likely to be accurate than the human’s judgment. Importantly, our findings reveal that AI advisors which do not account for the pillars can diminish value—even when using models that exceed human accuracy, while, interestingly, ReV-AI advisors can reliably add value to their contexts, even when the preferred model class cannot surpass human accuracy. Overall, the results demonstrate that the system-level, value-based approach we advocate for is essential to produce AI advisors that are dependable and can reliably benefit contexts, while highlighting the risks associated with alternative approaches.
How TeamRules’ distinct advising strategies benefit experts?
We now examine the distinct advising strategies that tr learns to reliably add value compared to alternative approaches and share insights on the role of the pillars and the selective advising mechanism. Towards this, we consider the Heart Disease diagnosis (classification) task [47], and a context inspired by evidence of physicians misdiagnosing heart failure in women at a higher rate as compared to men [56, 43, 57]. We simulate an expert with the reported performance bias who is also highly confident in their diagnoses for both men and women. While the expert would benefit from cost-effectively improving their decisions on women, it is challenging to produce advice that is inherently convincing to the expert given it ought to be even more confident than the expert’s (miscalibrated) overconfidence. The expert’s overconfidence thus limits the scope to improve their decisions cost-effectively, yet, such improvements would be consequential.
Figure 4 shows measures reflecting the advising strategies produced by the AI advisors and Figure 3 shows the outcomes when each AI advisor is paired with the expert to produce decisions. To characterize tr’s distinct advising strategy we contrast it with that of the task-only and tr-no(adb) advisors. We find that tr adds the most value by both improving more of the expert’s final decisions (Figure 3d,e,f) while achieving this through fewer engagements (Figure 3g,h,i). tr’s distinct advising strategy is shaped by each of the pillars, including the expert’s overconfident ADB, decision behavior, and context’s trade-off. Specifically, note that given the expert’s near perfect decision accuracy and high confidence for male patients, the scope for cost-effective advising on men is rather limited. We find that tr advises the expert highly selectively relative to the alternatives (Figure 4g,h,i). While most of the patients are males, tr’s advice focuses on women (65% of the advice), which is the group for which there are more opportunities to benefit the context and for which tr’s advice ultimately adds the most value (Figure 3a). As shown, tr’s conservative advising strategy on men yields a small improvement in decisions for men while avoiding the large losses from cost-ineffective human engagements incurred by the alternatives.
We also learn how the nature of tr’s advice is distinctly advantageous. tr advises with highly accurate (Figures 4d,e,f) and confident advice relative to alternatives (Figures 4a,b,c), leading the overconfident expert to be around twice as likely to accept tr’s advice relative to task-only’s advice (Figures 4j,k,i), and thereby improves more decisions cost-effectively. In contrast, the task-only advisor engages the human much more often with contradictory, yet, less accurate and less confident advice on both populations, ultimately harming value. Additionally, the tr-no(adb) advisor, which learns and selectively offers advice tailored to the human’s expertise and the context’s trade-off, and thereby provides highly accurate advice focusing on the female population, does not provide advice that is simultaneously sufficiently inherently convincing to the expert. Consequently, tr-no(adb)’s advice is accepted less often, and is thereby not cost-effective enough to warrant the human’s engagement.
In Appendix C.4, we discuss a complementary case study where the expert’s lower confidence makes them more receptive of advice for female patients. In this case, tr adapts to the human’s receptiveness by advising the expert on a higher proportion of female patients with lower-confidence advice that is sufficiently accurate and persuasive to cost-effectively improve more of the expert’s decisions..
3 Discussion
Although AI advisors can outperform human experts in some tasks, they risk diminishing value across AIaDM contexts. This can harm high-stakes practices, erode trust in AI, and slow progress. To enhance the positive impact of AI advisors, we advocate for a system-level, value-driven perspective. This shift is essential to understanding the failures observed in practical applications and to developing AI advisors that are reliable. Central to this approach are key pillars that characterize how AI advice can either add to or detract from value in practice. Our results establish that Rev-AI advisors can be relied on to add value across settings, and that neglecting any pillar risks diminishing value. The pillars also highlight the challenges of creating value-adding systems, such as understanding the human engagement costs, aligning with the context’s tolerance for those costs, and recognizing how the inherent properties of advice, besides its accuracy, influence its acceptance and added value.
This work thus provides critical insight into why AI advising systems struggle to add value in practice, and underscores that a system-level and value-based perspective is crucial to reliably add value in AIaDM contexts and enhance trust in AI systems. Our work provides timely and actionable guidance and a value-based framework to catalyze research on AI advising systems and foster their positive impact in practice.
4 Methods
The ReV-AI learning framework can be used to derive different concrete methods for generating ReV-AI advisors of different model classes. In this work, we develop and evaluate a rule-based prototype algorithm, TeamRules (tr), which produces ReV-AI advisors that generate decision recommendations in the form of inherently interpretable rules. The tr advising process during deployment is shown in Figure 1b. Inherently interpretable models are especially suitable for and are commonly used to advise humans in high-stakes contexts [31, 33, 34, 35, 36, 37, 38, 39] because they faithfully represent the model’s actual decision-making process [40].
4.1 Problem Formulation
We now formalize the framework for producing ReV-AI advisors shown in Figure 1a. Key notation for this section is summarized in Table 2. Let $\mathcal{X}$ be the set of decision instances with labels $\mathcal{Y}=\{0,1\}$, where $\mathcal{X}\times\mathcal{Y}\sim{\mathcal{D}}$. A decision instance $i$ is represented by features $x_{i}$, and outcome $y_{i}$. We consider contexts reflecting common high-risk decision domains in which a human first makes an independent decision $h_{i}\in\{0,1\}$. The human is then advised by an AI advising model, denoted by $m$, that either withholds advice or offers advice in the form of a decision recommendation $\hat{y}_{i}$. The AI recommendation is complemented by the AI model’s confidence [^5] We consider the model’s confidence because the impact of model confidence on human likelihood to accept advice has been studied in prior work. The framework can be instantiated with other forms of complementary ADB-bound information under the condition that the impact the ADB-Bound information has on ADB is modeled. $c^{M}_{i}$ reflecting the AI’s assessment of the likelihood that it’s advice is accurate. Finally, the human produces a final decision, by either accepting the recommendation $\hat{y}_{i}$ or selecting their own independent decision $h_{i}$.
When the AI does not offer advice, the human’s independent decision becomes the human-AI team’s final decision; when the AI offers advice that contradicts the human’s initial judgment, i.e., ${h}_{i}\neq\hat{y}_{i}$, the human applies their discretion before arriving at a final decision. We denote with $a_{i}$ whether the human DM accepts the AI recommendation for instance $i$, where $a_{i}=1$ denotes that the human DM accepts the AI advice, and $a_{i}=0$ means the DM rejects it. The human’s discretion over AI advice is represented by the likelihood of the human accepting the AI advice $p(a|c^{M},c^{H},\hat{y}\neq h)$, and is estimated with $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)$, where $c^{H}$ represents the human DM’s confidence in their original decision $h$. Throughout the rest of the paper, we refer to $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)$ as the human’s estimated discretion model and denote it with $\hat{p}(a)$, for brevity.
Table 2: Key Notation
Notation Description
$x_{i}$ Feature values for instance at index i.
$y_{i}$ Ground truth label for instance at index i.
$h_{i}$ Human expert’s initial decision (before seeing AI advice) for instance at index i.
$c^{H}_{i}$ Human expert’s confidence in their initial decision for instance at index i.
$c^{M}_{i}$ Model’s reported confidence in its decision advice for instance at index i.
$\hat{y}_{i}$ Model’s decision advice for instance i.
$\hat{p}(a|c^{M},c^{H},h\neq\hat{y}))$, denoted with $\hat{p}(a)$ for brevity Estimated model of human’s algorithm discretion behavior (ADB) - likelihood human accepts a contradicting recommendation given their confidence in their initial decision and model’s confidence in its advice.
$\hat{p}(y|x)$ Estimated model of ground truth label for an instance.
$R,R^{+},R^{-}$ Sets of rules. $R^{+}$ and $R^{-}$ denote sets of rules that correspond to a model producing positive and negative decision advice if covered by each rule set, respectively.
$\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)$ Expected team loss over human’s likelihood to accept a piece of advice.
$\mathcal{V}(y,d)$ Denotes the context-relevant loss incurred from the human’s final decision $d$ given ground truth $y$.
$\alpha$ Represents the context’s advising cost-benefit trade-off.
$\mathbbm{C}(R,x_{i})$ Rule coverage indicator. True if instance with feature values $x_{i}$ is covered by any rule in $R$.
$\psi(y_{i},y^{*}_{i},h_{i},\hat{p}(a),\alpha)$ Indicator function that determines whether a prospective advice $y^{*}$ adds value to the expert in expectation.
$\mathbbm{1}\{\text{condition}\}$ Indicator function. Returns $1$ if condition is met, 00 otherwise.
TTL Total team loss from final decisions.
ReV-AI’s advice $\hat{y}$ for a decision instance is either a decision recommendation $z\in\mathcal{Y}$ or the lack of a decision recommendation, such that the withholding of advice corresponds to recommending the human’s initial judgment $h$. Formally, $\hat{y}$ is given by:
$\hat{y}$ =$=\begin{cases}z\in\mathcal{Y},&\text{if decision advice is offered}\\ h,&\text{if advice is withheld}\end{cases}$ (1)
Simultaneously, the selective advising decision – whether to provide advice and the specific advice $z$ – is personalized to the human and context’s cost-benefit trade-off, with the goal of maximizing the advisor’s added value:
$m^{*}=\underset{m\in\mathcal{M}}{\operatorname*{arg\,min}}\bigg{[}\mathbb{E}_{x,y,h\sim\mathcal{D}}\bigg{(}\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)\bigg{)}\bigg{]}$ (2)
where $\mathbb{E}_{x,y,h\sim\mathcal{D}}$ denotes expectation over the input space $(X,Y)$ and the human’s independent decisions over it with joint distribution $D$; $\mathcal{L}$ is the expected team loss over the human’s discretion $a$ resulting from the given AI advice, the given instance $(x,y)$, and and the human’s independent decision $h$, given by:
$\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)$ $=\hat{p}(a)\mathcal{V}\big{(}y,\hat{y}\big{)}$ (3)
$\text{\indent}+\big{(}1-\hat{p}(a)\big{)}\mathcal{V}(y,h)+\alpha\mathbbm{1}\{\hat{y}\neq h\}$
where $\mathcal{V}(y,d)$ denotes the context-relevant loss incurred from the human’s final decision $d$ given ground truth $y$, and $\alpha$ denotes the context’s trade-off between reconciling contradictory advice and benefits from improving decisions, such that the loss from reconciling one contradictory advice is equivalent to the following loss from a decision error: $\frac{\alpha}{\mathcal{V}(y,d|y\neq d)}$. Equivalently, the trade-off implies that $\frac{\mathcal{V}(y,d|y\neq d)}{\alpha}$ contradicting pieces of advice produce loss equivalent to that of a single incorrect decision. Note that Equation (3) reflects that when $m$ offers contradictory advice, i.e., when $\hat{y}\neq h$, the expected loss depends on the estimated likelihood $\hat{p}(a)$ that the human accepts the advice; the loss in this case also includes the cost $\alpha$ from the human’s reconciliation of the contradiction. When $m$ does not provide decision advice or when it provides advice that agrees with the human’s initial judgment $h$, then $\hat{y}=h$, and the loss becomes $\mathcal{V}(y,h)$.
The total expected loss produced by an advisor, $\mathbb{E}_{x,y,h\sim\mathcal{D}}\bigg{(}\mathcal{L}(y,\hat{y},h,\hat{p}(a),\alpha)\bigg{)}$, is henceforth referred to as the Total Team Loss (TTL). Importantly, as reflected in our discussion and shown in Equation (3), the loss includes the loss from reconciling contradictions in addition to the loss from incorrect final decisions. Ultimately, offering value-adding advice hinges on improvements achieved cost-effectively per the context’s cost-benefit trade-off.
The ReV-AI advisor is informed by the underlying data domain $\mathcal{X}\times\mathcal{Y}$, data reflecting the human’s expertise, confidence, and ADB for estimating $p(a|c^{M},c^{H},\hat{y}\neq h)$, and the context’s tolerance for added costs to improve decisions. Data reflecting the human’s expertise - consisting of the human’s independent decisions =$\mathcal{H}=\{0,1\}$ for instances in $\mathcal{X}$ is often available in historical data and which otherwise can be acquired for the corresponding domain. Prior work also established how a given individual’s likelihood of accepting a given piece of AI advice for a given decision, $\hat{p}(a)$, can be effectively estimated based on the human’s self-reported (possibly miscalibrated) confidence in their decision and the AI’s confidence, $c^{H}$ and $c^{M}$, respectively [27, 28]. Following prior work, for each instance $(x,y)$ for which advice is offered, the human’s decision $h$, self-reported confidence $c^{H}$, and decision to accept or reject the advice $a$, is acquired in a risk-free setting, prior to deployment. Prior to producing an initial ReV-AI advisor, data can be acquired for advice offered by a general-purpose model trained on $\mathcal{X}\times\mathcal{Y}$, based on which $\hat{p}(a)$ can be modeled and an initial ReV-AI can be produced. Importantly, data on both human expertise and discretion can be acquired periodically to reflect changes. We discuss modeling the human’s discretion in more detail in Appendix A.2 and we study the robustness of the approach to varying degrees of accuracy for these predictions in Appendix C.5.
4.2 Producing TeamRules
Our framework can be used to derive different methods that generate ReV-AI advisors for different humans and contexts. We develop a rule-based prototype advisor, TeamRules (tr). Rules have been shown to offer an especially intuitive form of inherently interpretable advice for high-stakes decision-makers, given their transparent inner structures and expressiveness [31, 33, 34, 35, 36]. The purpose of the prototype is to first demonstrate the use of the framework to derive a method for generating ReV-AI advisors. Most importantly, the prototype allows us to study and characterize how the behaviors of different ReV-AI advisors’ produced by (tr) vary across contexts, and to study the ReV-AI advisors’ impact on value added to the human relative to that of alternatives.
We begin by formalizing the terms rule and rule set used to advise a decision-maker:
Definition 1 (Rule).
A rule $r$ is a logical expression made of conditional statements about a conjunction of a subset of feature values. $r$ covers an example $x_{i}$, denoted as $\mathbbm{C}(r,x_{i})=1$, if the logical expression evaluates to true given $x_{i}$.
Definition 2 (Rule Set).
A rule set $R$ is a collection of rules. $R$ covers an example $x_{i}$, denoted $\mathbbm{C}(R,x_{i})=\mathbbm{1}\{\sum_{r\in R}\mathbbm{C}(r,x_{i})\geq 1\}$, if at least one rule $r$ in $R$ covers $x_{i}$.
Given a training set of instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, tr learns a set of rules $R=\{R^{+},R^{-}\}$ to advise the human, such that rules in $R^{+}$ or $R^{-}$ can be used to advise 1 or 0 on instances that satisfy them, respectively. Note that a distinct property of rule-based advice is that a given rule set $R$ can be produced so that it does not cover all instances. Following our framework, to minimize the expected loss from the human’s final decisions, a tr advisor selectively offers advice for an instance $x_{i}$ if $x_{i}$ is covered by a rule in $R$ (i.e., if $\mathbbm{C}(R,x_{i})=1$) and if the expected Total Team Loss (TTL) from offering the advice is lower than the expected TTL when advice is withheld. In summary, Figure 1b demonstrates TR’s advising process which consists first of providing possible advice determined by the rule set $R$ and by further assessing the expected added value of advice when an instance is covered by $R$.
The learned rule set $R$ simultaneously determines (a) for which decision instances advice may be offered, and (b)what rule-based advice can be offered to the human so that it minimizes the TTL. Below we outline the optimization method for deriving $R$ that minimizes the TTL from the human’s final decisions by accounting for the human’s expertise, ADB, and the context’s cost-benefit trade-off.
Selective Advising. Formally, tr’s selective advising at inference time is outlined in Equations (4) and (5). For any decision instance $x_{i}$ encountered during deployment, tr either offers advice or withholds advice which is equivalent to recommending the human’s independent decision $h_{i}$, given by: [^6] Note how Equation 1 from the ReV-AI advising framework is concretely operationalized to form tr’s rule-based advising process shown in Equation 4.
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbbm{C}(R^{+},x_{i})\wedge\psi(y,1,h_{i},\hat{p}(a),\alpha)\\ 0,&\text{if }\mathbbm{C}(R^{-},x_{i})\wedge\neg\mathbbm{C}(R^{+},x_{i})\wedge\psi(y,0,h_{i},\hat{p}(a),\alpha)\\ h_{i},&\text{otherwise}\end{cases}$ (4)
Where $\psi(y,y^{*},h,\hat{p}(a),\alpha)$ shown below is the indicator function that determines whether a prospective advice $y^{*}$ adds value in expectation, i.e., whether the expected loss from the advice ($\mathbbm{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$) is lower than the expected loss from withholding advice ($\mathbbm{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]$:
$\psi(y_{i},y^{*}_{i},h_{i},\hat{p}(a),\alpha)$ $=\mathbbm{1}\{\mathbbm{E}_{y}[\mathcal{L}(y,y^{*}_{i},h_{i},\hat{p}(a),\alpha)]$ (5)
$\text{\indent}<\mathbbm{E}_{y}[\mathcal{L}(y,h_{i},h_{i},\hat{p}(a),\alpha)]\}$
To take expectation over the ground truth labels, $\mathbbm{E}_{y}$, we train a model $\hat{p}(y|x)$ using an XGBoost classifier [58] with hyperparameters selected using grid-search cross-validation [59]. In addition, the team loss from the human’s final decision for instance $x_{i}$, $\mathcal{L}(y_{i},\hat{y_{i}},h_{i},\hat{p}(a),\alpha)$, is shown in Equation (3). In our implementation of tr, we set the instance-level decision loss in Equation (3) to: $\mathcal{V}(y,d)=\lambda_{0}^{(1-y)}\lambda_{1}^{(y)}\mathbbm{1}(y\neq d)$, where $\lambda_{0}$ and $\lambda_{1}$ are the decision losses incurred from false positives and false negatives, respectively. All main results consider settings in which $\lambda_{0}=\lambda_{1}=1$, however, we also include additional results where $\lambda_{0}\neq\lambda_{1}$ in the Appendix.
tr accompanies advice with ADB-bound information, namely, its corresponding confidence $c^{M}_{i}$, which is the precision of the rule used to produce the advice. When an instance is covered by multiple rules producing the same outcome, the highest precision rule is used. Specifically, the confidence is given by:
$c^{M}_{i}$ $=\underset{r\in R^{sign(\hat{y}_{i})}:\mathbbm{C}(r,x_{i})=1}{\max}\Bigg{[}\frac{\sum_{j=1}^{n}\big{(}\mathbbm{1}\{\mathbbm{C}(r,x_{j})\wedge\mathbbm{1}(y_{j}=\hat{y}_{i})\}\big{)}}{\sum_{k=1}^{n}\mathbbm{C}(r,x_{k})}\Bigg{]}$ (6)
Note that the confidence provided is an inherent characterization of the quality of the rule used to provide the piece of advice.
Algorithm 1 TR Rules Generator
1:Input: Data $(x_{i},h_{i},c^{H}_{i},y_{i})\text{ }\forall i\in\{0,...,n\}$, Discretion Model $\hat{p}(a)$
2:Parameters: $T,\alpha,C_{0},\beta_{0},\beta_{1},\beta_{2}$
[…]
5:$R_{0}\leftarrow\emptyset$, $R^{*}\leftarrow R_{0}$
6:for $t=1...T$ do
7:     $R_{t}\leftarrow R_{t-1}$
8:     for $i=1...n$ do
9:         Produce $\text{Produce }\hat{y}_{i}$, $c^{M}_{i}$triangleright According to Equations (4) and (6).
10:     ϵitalic-$\epsilon\leftarrow$ draw instance at index $i$ with probability: $\frac{\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\alpha\mathbbm{1}\{\mathcal{V}(y_{i},1-y_{i})\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})>\alpha\}}{\sum_{j=1}^{n}\bigg{(}\mathcal{L}(y_{j},\hat{y}_{j},h_{j},\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j}),\alpha)-\alpha\mathbbm{1}\{\mathcal{V}(y_{j},1-y_{j})\hat{p}(a_{j}|c^{M}_{j},c^{H}_{j},\hat{y}_{j}\neq h_{j})>\alpha\}\bigg{)}}$
11:     if ($\mathbbm{C}(R_{t-1},x_{\epsilon})$ and ($\hat{y}_{\epsilon}\neq y_{\epsilon}$)) or (($h_{\epsilon}\neq\hat{y}_{\epsilon}$) and ($\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})<\alpha$)) then
12:         if $(y_{\epsilon}=0)$ then
13:              $R_{t}\leftarrow$ cut or replace rule from $R^{+}_{t}$ that covers $x_{\epsilon}$
14:         else
15:              $R_{t}\leftarrow$ cut or replace rule from $R^{-}_{t}$ that covers $x_{\epsilon}$
16:     else
17:         $R_{t}\leftarrow$ add rule to $R^{sign(y_{\epsilon})}_{t}$
18:     for $i=1...n$ do
19:         Produce $\text{Produce }\hat{y}^{*}_{i}$, $c^{*M}_{i}$triangleright Using updated rule-set $R_{t}$
20:     if $\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}_{i}^{*},h_{i},\hat{p}(a_{i}|c^{*M}_{i},c^{H}_{i},\hat{y}_{i}^{*}\neq h_{i}),\alpha)<\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)$ then
21:         $R^{*}\leftarrow R_{t}$
22:     if $\exp\left({\frac{\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}_{i},h_{i},\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i}),\alpha)-\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y}^{*}_{i},h_{i},\hat{p}(a_{i}|c^{*M}_{i},c^{H}_{i},\hat{y}^{*}_{i}\neq h_{i}),\alpha)}{C_{0}^{\frac{t}{T}}}}\right)\leq$ random$([0,1])$ then
23:         $R_{t}\leftarrow R_{t-1}$
24:Output: $R^{*}$
4.3 Optimization Approach
We develop a training algorithm based on the core idea of simulated annealing to learn a rule set $R$ that minimizes the team’s loss in Equation (3). Simulated annealing is well-suited for our discrete space, and has been used successfully in previous rule-based models [60, 61]. The complete algorithm is shown in Algorithm 1. The optimization procedure takes as input the set of training instances $D=\{x_{i},y_{i},h_{i},c^{H}_{i}\}_{1}^{n}$, along with the cost of reconciling a piece of contradictory advice $\alpha$ and the human discretion model $\hat{p}(a)$. Initially, a set $\Gamma$ of $\beta_{0}$ candidate rules is generated using an off-the-shelf rule mining algorithm, such as FP-growth [62]. We follow prior work and established practices to set the hyperparameters which we discuss in Appendix A.3. The initial rule set $R_{o}$ is set to an empty set so that decisions are made exclusively by the human. At each subsequent iteration $t$, a new rule set is proposed by either adding or dropping a rule to the current solution. This new set is accepted with a probability that depends on the quality of the proposed solution — i.e., the lower the total team loss (TTL) across all training instances, the higher the likelihood of acceptance. The probability also depends on the temperature, which decreases over iterations — i.e., the higher the temperature, the greater the likelihood of acceptance. As a result, the algorithm explores more aggressively at the beginning of the search, being more likely to accept solutions even if they are worse than the current one to thoroughly explore the space. Towards the end, it becomes more conservative, accepting only solutions that improve the existing rule set. This approach allows the algorithm to progressively minimize the TTL while avoiding getting stuck in local minima.
By following the principle of stochastic local search [63], it is possible to explore changes to R by either adding to or removing from $R$ a randomly selected rule. However, to expedite convergence, revisions to $R$ are guided by identifying instances that contribute to the loss and exploring changes that may decrease this loss. Specifically, at each iteration, an instance with index $\epsilon$ is drawn at random with the probability of being drawn proportional to its relative contribution to the empirical TTL over all instances, as shown in Line 9 Algorithm 1. Note, however, that the contradiction loss is ignored from the loss contribution if the possible improvement in decision loss $\mathcal{V}(y_{i},1-y_{i})\hat{p}(a_{i}|c^{M}_{i},c^{H}_{i},\hat{y}_{i}\neq h_{i})$ is greater than the incurred reconciliation cost $\alpha$. We do not account for the reconciliation cost in such cases because the benefit from attempting to change the human’s decision is greater than the cost incurred from reconciling, so we do not treat such examples as needing a correction, despite them contributing to the final TTL. Once an instance is selected, $R_{t}$ may be adapted by either adding, removing, or replacing a rule if the adaptation decreases the empirical TTL. Subsequently, the change is accepted only if it reduces the empirical TTL, i.e., if $\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y^{*}}_{i},h_{i},\hat{p}(a),\alpha)<\sum_{i=1}^{n}\mathcal{L}(y_{i},\hat{y_{i}},h_{i},\hat{p}(a),\alpha)$. Finally, to encourage exploration, we reset $R_{t}\leftarrow R_{t-1}$ even if the loss produced by $R_{t}$ is higher than the loss produced by $R_{t-1}$ with a decreasing probability over time, as shown in Line 21, Alg. 1 and as is typically done in simulated annealing procedures [62].
Appendix A Experiment Details
Here we provide full details on datasets and simulation procedures used for all experiments.
A.1 Heart Disease, FICO, HR Datasets
Instance counts for all datasets are shown in Table 3.
Table 3: Dataset Instance Counts
Dataset Train Instances Validation Instances Test Instances
Heart Disease 505 87 127
FICO 6,120 801 1,080
HR 568 38 143
A.2 Human Behavior Simulation Details
These details correspond to the results shown in Figure 2.
A.2.1 Step 1: Simulating Human Decisions
Following prior work which simulated human decision-behavior for human-AI collaborative settings [49, 55, 53], we represent human decision behavior as the probability of the human correctly determining the label $y_{i}$ for any instance $i$. Specifically, we create two types of decision behaviors: (1) Difficulty-Biased and (2) Group-Biased.
Difficulty-Biased behavior is simulated by first training an out-of-the-box probabilistic model, such as logistic regression, to predict the correct label for unseen task instances. We can denote the probabilistic output of the model as $\hat{p}_{logistic}(y|x)$, and we use this output as a proxy for the difficulty of the instance. Specifically, we denote the proxy value for instance $i$’s difficulty with $d_{i}$, such that $d_{i}=2|\hat{p}_{logistic}(y=1|x)-0.5|$. If $d_{i}>d_{j}$, then we consider instance $i$ as easier to classify than instance $j$, because the probabilistic model predicts its class with a higher likelihood (is more confident of the instance’s label). Finally, we partition the instance space such that each value $d_{i}$ for all instances $i$ maps to the simulated human’s likelihood of being correct, and we create regions such that the human has a high accuracy for some difficult instances, and a low accuracy for some easier instances. Specifically, we set the human’s accuracy to $60\%$ whenever $d>d_{t}$ and $100\%$ whenever $d<=d_{t}$, where $d_{t}$ is a dataset specific difficulty threshold. $d_{t}$ is set to $0.6$ for the Heart Disease dataset, $0.3$ for the FICO dataset, and $0.8$ for the HR dataset. Values were chosen so that the human would have $100\%$ accuracy on a smaller part of the dataset (roughly 25%-30% of instances) and 60% on the rest, leading each human to have an overall accuracy ranging from 70%-80%.
Group-Biased behavior is simulated by selecting a feature upon which to partition the instance-space. For example, in practice, some doctors have been shown to exhibit gender and age bias when making heart disease diagnostic decisions [42]. Simulating group-biased behavior thus involves selecting a feature that partitions the space into groups, such as gender, and subsequently setting the simulated DM’s likelihoods of being correct to different values for each partition. Such a partition allows us to simulate expert decision-making biases often seen in practice. For each dataset, if an instance meets a feature-based condition, the human’s accuracy is set to 60%, otherwise, it is set to 95%. The conditions for each dataset are as follows:
  • Heart Disease: Age $<50$
  • FICO: Number of Satisfactory Trades $<24$
  • HR: Age $>32$
A.2.2 Step 2: Simulating Human Confidence in Their Decisions
Next, given a simulated human decision-behavior, we simulate human confidence behaviors. A decision-maker (DM) with accuracy-biased confidence exhibits low confidence in decisions they are less likely to get correct and high confidence on decisions they are more likely to get correct. Specifically, the accuracy-biased human has confidence $c^{H}_{i}=p(h_{i}=y_{i})\pm\kappa$, meaning their confidence equals the likelihood of making a correct decision, offset by a constant value. This offset term ensures that the human’s probability of being correct on a given instance is not exactly equal to their confidence, as humans rarely have perfectly calibrated confidence in their decisions [44, 45, 28]. We also generate group-biased DM confidence behavior. Humans with group-biased confidence behavior have confidence that varies with both instance difficulty and along a selected feature. For instance difficulty, we use the same difficulty values and thresholds $d_{t}$ described in Section A.2.1. The features selected for each dataset are gender for the Heart Disease and HR datasets, and external risk estimate score for the FICO dataset. For all datasets, when $d<d_{t}$, confidence is set to $0.9$. When $d>=d_{t}$, we use the following group-based rules:
  • Heart Disease & HR: Confidence is set to $0.2$ when gender is male, otherwise confidence set to $1$.
  • FICO: Confidence is set to $0.2$ when external risk estimate $<65$, otherwise confidence set to $1$.
A.2.3 Step 3: Simulating and Modeling Human ADB
To simulate humans’ ground truth ADB, $p(a|c^{M},c^{H},\hat{y}\neq h)$, e.g., the likelihood with which they would accept contradicting AI advice, we follow prior work that explored various model forms of humans’ ADB as functions of their self-reported (and possibly miscalibrated) confidences in their decisions and the AI advisor’s reported confidence values [27]. Specifically, we utilize the Adjusted Naïve Bayes (ANB) CPT-Based utility component:
$\hat{u}^{t}_{accept}=(1+\beta_{ADB})w(\frac{1}{1+\frac{(1-a)\dot{(}a-b)}{a\dot{b}}})-\beta_{ADB}$ (7)
$\hat{u}^{t}_{reject}=1-(1+\beta_{ADB})w(\frac{1}{1+\frac{(1-a)\dot{(}a-b)}{a\dot{b}}})$ (8)
Where:
$a=\frac{(c^{M})^{\gamma}}{(c^{M})^{\gamma}+(1-c^{M})^{\gamma}}$ (9)
$b=\frac{(c^{H})^{\gamma}}{(c^{H})^{\gamma}+(1-c^{H})^{\gamma}}$ (10)
$w(p)=\frac{(p^{k})}{p^{k}+(1-p)^{k}}$ (11)
The ANB CPT-Based utility component characterizes how the human DM estimates the utility of accepting and rejecting AI advice. We then utilize the basic Logit selection component:
$p(a|c^{M},c^{H},\hat{y}\neq h)=\frac{\exp(\delta*\hat{u}^{t}_{accept})}{\exp(\delta*\hat{u}^{t}_{accept})+\exp(\delta*\hat{u}^{t}_{reject})}$ (12)
The selection component characterizes how the human DM stochastically decides whether to accept AI advice given their estimate of the utility provided from accepting AI advice. We selected the basic Logit selection component because applying human-based adjustment on the utility component of the two-component model always helps to increase the model’s predictive performance if the basic selection component was Logit [27]. Additionally, utilizing the basic Logit component rather than the human-adjusted component allows us to make fewer model parameter assumptions. We select the ANB CPT-Based utility component because it was found to generally lead to models that best fit to human behavior across treatments and regardless of selection component choice [27].
We utilize the following parameter values for our simulation of human ADB: $\delta=5$, $k=0.63$, $\gamma=0.95$, $\beta_{ADB}=0.5$. The $\delta$ parameter model reflects the human decision maker’s sensitivity to their estimates of utility gained/lost from correct/incorrect decisions: when $\delta\rightarrow 0$, the human decision maker takes actions randomly; when $\delta\rightarrow\infty$, the human decision maker always takes the action with the optimal estimated utility. We select a value such that the human would generally take the action with optimal estimated utility but not always. The human determines their estimated utility from correct/incorrect decisions by transforming their confidence and the AI’s reported confidence through several functions that rely on parameters $k,\gamma$, and $\beta_{ADB}$.
The parameter $\gamma$ varies from 0 to 1 and controls the degree to which the human adjusts their own confidence values and the AI’s reported confidence values to $0.5$ before integrating them into a single value that represents the human’s confidence in the AI’s decision. When $\gamma=0$, all the probabilities are transformed to 0.5, while when $\gamma=1$, no adjustment is applied to probabilities before the human integrates the confidences. We select $\gamma=0.95$ to represent a human that slightly adjusts probabilities closer to 0.5. Once the human has an estimate of the AI’s probability of being correct, this probability is further transformed into a perceived probability, motivated by human tendency to perceive probabilities non-linearly [27]. The $k$ parameter determines how the likelihood they believe the AI to be correct is transformed into a perceived probability. We select a value of $k=0.63$, which corresponds to an inverse-S shaped probability weighting function, i.e. the human overweights extreme events, and underweights common events. We select this value because expert DM’s tend to overestimate the likelihood of rare diseases (rare event) and thus underestimate the lack of rare diseases (common event) [64]. Finally, given the human’s final perceived probability of the AI being correct, $w(p)$, the $\beta_{ADB}$ parameter determines how much utility they expect to gain from accepting/rejecting AI advice. The $\beta_{ADB}$ parameter can be used to increase the disparity between utility gained from decisions expected to be correct and incorrect. A $\beta_{ADB}$ value of 0 means the expected utility gained from accepting advice is equal to $w(p)$ and the expected utility gained from rejecting advice is equal to $1-w(p)$. As $\beta_{ADB}$ increases, the gap between expected utility gains increases, corresponding to higher stakes settings. We thus set $\beta_{ADB}$ to 0.5, slightly increasing the perceived stakes across all settings in our main results.
Recall that, given the human’s ADB is unknown, the ReV-AI learning framework requires the collection of data and subsequently induces a model to estimate the human’s likelihood of accepting AI’s advice for any given instance: $p(a|c_{M},c_{H},\hat{y}\neq h)$. Following our framework, for a subset of decision instances we generate data reflecting the simulated human’s independent decisions, confidence, and ADB in response to an AI recommendation and confidence. The AI recommendation and confidence is provided to the human initially by a logistic regression model trained on a withheld subset of data. In the experiments reported here, we then induce $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)$ from the interaction data between the human and the logistic regression model using an out-of-the-box implementation of XGBoost [58].
A.3 Training Parameters
  • $T=2000$
  • $C_{0}=0.01$
  • $\beta_{0}=0.05$
  • $\beta_{1}=4$
  • $\beta_{2}=10,000$
We follow prior work and established practices to set the maximum rule length $\beta_{1}$ and minimum support $\beta_{2}$ [61, 65, 66]. Specifically, rule-based models utilize rules of lengths 2-4 and a minimum support threshold $\beta_{2}$ of 5%. These broadly accepted standards are informed by consistent findings for settings that reduce the risk of overfitting and and promote intepretability.
A.4 Measures of Performance.
[…]
Appendix B Alternative Rule-Production Methods
Task-Only
The task-only method, inspired by the Bayesian Rule Sets method [61], produces a rule set that is optimized for maximum generalization accuracy on the task data and which covers all instances (provides advice on all instances). Accordingly, the task-only approach offers decision advice $\hat{y}_{i}$ given by:
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbbm{C}(R,x_{i})\\ 0,&\text{otherwise}\end{cases}$ (13)
And it offers model confidence $c^{M}_{i}$:
$c^{M}_{i}$ =$=\begin{cases}\underset{r\in R:\mathbbm{C}(r,x_{i})=1}{\max}\Bigg{[}\frac{\sum_{j=1}^{n}\big{(}\mathbbm{1}\{\mathbbm{C}(r,x_{j})\wedge\mathbbm{1}(y_{j}=\hat{y}_{i})\}\big{)}}{\sum_{k=1}^{n}\mathbbm{C}(r,x_{k})}\Bigg{]},&\text{ if }\hat{y}_{i}=1\\ \frac{\sum_{j=1}^{n}\mathbbm{1}\{y_{j}=\hat{y}_{i}\wedge y_{j}=0\}}{\sum_{k=1}^{n}\mathbbm{1}(\hat{y}_{i}=0)},&\text{ otherwise }\end{cases}$ (14)
The loss is given by:
$\mathcal{L}_{task-only}(y,\hat{y})=\mathcal{V}(y,\hat{y})$
where $\mathcal{V}(y,d)$ is the same context relevant loss utilized in the corresponding tr model. The full optimization algorithm used for the task-only approach is Algorithm 1 from [61].
TR-no(ADB, Cost), TR-no(ADB), TR-no(Cost)
The tr-no(adb, cost), tr-no(adb), and tr-no(cost) frameworks produce rule sets and provide advice for some instances, but may defer to the human if their advice is not expected to add value given the pillars they consider. The tr-no(adb, cost) approach accounts for human expertise, the tr-no(adb) approach accounts for both human expertise and the context’s advising cost-benefit trade-off, and the tr-no(cost) approach accounts for both human expertise and ADB. All approaches are variants of tr and work similarly. Models that do not account for ADB use a fixed discretion model $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)$ as input. Whereas $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)$ is an estimate of the human’s ground truth discretion behavior in tr, $\hat{p}(a|c^{M},c^{H},\hat{y}\neq h)=1$ in tr-no(ADB) and tr-no(adb, cost), meaning that these models are trained with the assumption that the human would accept all advice offered (does not consider their ADB). Next, the methods that do not account for the cost-benefit trade-off, tr-no(adb, cost) and tr-no(cost) are optimized with $\alpha=0$, mean they assume there is no cost to providing AI advice, regardless of the context.
Figure 5: Value added by AI advisors relative to human for environments with varying yet excessively inhibitive trade-offs between advising costs and decision benefits for Heart Disease, FICO, and HR datasets. X axis reflects contexts in which the cost from 1/x human engagements are equivalent to the loss from one incorrect decision: higher x-values reflect that fewer human engagements are permissible to achieve a given decision benefit. Results show average value-added +- SE (shaded region) over 10 repetitions.
Appendix C Additional Results
In this section, we provide results on additional settings and offer additional analysis on the benefits offered by tr’s mechanisms. Specifically, we demonstrate how tr can add more value than alternatives when advising costs are especially inhibitive ($\alpha>0.5$) and when the decision loss incurred from false positives and false negatives is unequal. We provide an additional case study to assess how tr adds value relative to alternatives when an expert decision-maker has confidence that is calibrated to their decision-making strengths and weaknesses. We also assess the value of personalizing the rules of a rule-based advisor - as is done in tr’s optimization scheme - by comparing to a version of tr in which the rules are optimized only for generalization accuracy on the task, but the selective advising mechanism is subsequently applied to each piece of advice generated by those rules. Finally, we also demonstrate how the value-added by tr decreases as discretion model accuracy decreases, and we demonstrate how a simple robustness mechanism can ensure that advisors optimized using low quality discretion models will not diminish the value offered by the human alone.
C.1 Prohibitive Advising Costs
In our main results, we show the value-added by tr relative to alternatives for $0<\alpha\leq 0.5$. In this section, Figure 5, we show how tr can also add some value and does not diminish the value offered by the human even when the advising cost-benefit trade-off is especially inhibitive ($0.6\leq\alpha\leq 1$). The alternative frameworks produce advisors that greatly diminish the human’s standalone value across settings, while tr generally learns that it cannot add any value and refrains from advising the human. In some cases, it is even able to add some marginal value, such as in some settings when $\alpha=0.6$.
C.2 Asymmetric Decision Loss
Often in high-stakes settings the loss incurred from false positives and false negatives is not equal. For example, in disease diagnostics, failing to correctly diagnose a patient with a disease is a more costly error than diagnosing that a patient has a disease when they do not. We consider an additional setting for the heart disease classification task, one in which $(\lambda_{0},\lambda_{1})=(1,3)$. This setting corresponds to false negatives contributing three times as much error as false positives, respectively. We consider an expert that exhibits a group-based decision bias; The expert has a high accuracy ($95\%$) on the young and female populations, and a low accuracy ($60\%$) on the elderly-male population. We find that tr significantly improves the human’s overall decisions more than the task-only and even the tr-no(adb) approaches in this setting, shown by the lower TTL from the Total population in Table 4. Note that all methods are optimized to account for the asymmetric decision loss. TR leads to the greatest reduction in false negatives while also leading to the lowest TTL. This result demonstrates yet another practical context in which ReV-AI advisors can add superior value over alternatives.
Table 4: Advising outcomes in terms empirical Total Team Loss (TTL) $\pm$ SE after 20 repetitions on the heart disease diagnosis task in the presence of asymmetric costs from false negatives (cost $3$) and positives (cost $1$) and when there are no reconciliation costs incurred. We pair the tr, tr-no(adb), and Task-Only advisor with a simulated individual DM. We include the tr-no(ADB) advisor to demonstrate value of leveraging ADB in high-risk, asymmetric cost settings, even when there are no reconciliation costs incurred by the DM.
TR TR-No(ADB) Task-Only Human-Alone
No Disease 0.116 $\pm$ 0.005 0.099 $\pm$ 0.003 0.150 $\pm$ 0.005 0.099 $\pm$ 0.003
Has Disease 0.115 $\pm$ 0.005 0.145 $\pm$ 0.008 0.156 $\pm$ 0.012 0.213 $\pm$ 0.007
Total 0.231 $\pm$ 0.006 0.243 $\pm$ 0.005 0.307 $\pm$ 0.015 0.313 $\pm$ 0.007
C.3 Personalizing Advising Rules
In Figure 6, we compare the performance of the tr advisors to tr-SelectiveOnly advisors across contexts. The tr-SelectiveOnly advisors are not optimized to produce value-maximizing advice, instead, they are optimized equivalently to the task-only approach. However, we apply tr’s selective advising mechanism to only offer the advice from tr-SelectiveOnly if it is expected to add value given all the pillars. Therefore, the tr-SelectiveOnly advisor will only offer advice if it is expected to add value, but what advice it can offer is not personalized to add value to the human. We find that tr can add significantly more value than tr-SelectiveOnly in some settings (Figures 6b,g,i), demonstrating the potential benefit of personalizing the rules such that what advice the model can give is sufficiently accurate and convincing to add value to the given expert. Notably, in other settings, tr leads to similar value-added compared to tr-SelectiveOnly. This demonstrates that personalizing advice to add value does not diminish value across settings, and may add value in some settings.
Figure 6: Value added by AI advisors relative to human for environments with varying trade-offs between advising costs and decision benefits for Heart Disease, FICO, and HR datasets. X axis reflects contexts in which the cost from 1/x human engagements are equivalent to the loss from one incorrect decision: higher x-values reflect that fewer human engagements are permissible to achieve a given decision benefit. Results show average value-added +- SE (shaded region) over 10 repetitions.
C.4 Additional Case Study
Figure 7: Case Study - Outcomes: Advising outcomes when tr, task-only, tr-no(adb) advisors advise an expert with relatively appropriate confidence given their expertise. The context’s trade-off is $0.1$, reflecting that reconciling up to 10 contradictory pieces of AI advice is deemed cost-effective if the contradictions yield at least one improved decision. The expert has an independent decision accuracy of 90% on the Male population (majority), a 60% accuracy on the Female population (minority), and an overall accuracy of 87.5% on the entire population. This expert is highly confident (97.5% confidence) on the the male population and less confident (60.5% confidence) on the female population. Results are averages over 20 repetitions $\pm$ SE (vertical line at center of each bar).
Figure 8: Case Study - Advising Strategies: Properties of advice corresponding to Figure 7. Advising an expert with relatively appropriate confidence in their abilities leads TR to learn and selectively offer high accuracy advice that is sufficiently convincing given the expert’s confidence. Results from 20 repetitions.
We simulate an expert, who we will refer to as Doctor B and who exhibits the same decision behavior as the expert in the main results section, who we will refer to as Doctor A. Doctor B has high accuracy when diagnosing male patients, and low accuracy when diagnosing female patients. However, in contrast to Doctor A, Doctor B is aware of their weakness when diagnosing female patients, and thus has appropriately low confidence (60%) when diagnosing female patients while remaining slightly overconfident on male patients. Results are shown in Figures 7 and Figures 8.
We find that tr again focuses on providing majority of its advice for the female population, however, it adapts its advising strategy by producing lower confidence advice on the female population than when advising Doctor A. This allows for tr to offer more advice to Doctor B than to Doctor A, and because the threshold for providing cost-effective advice is lowered given Doctor B’s low confidence and consequential receptiveness to advice, more decisions can be cost-effectively improved. tr also produces a different advising strategy when advising Doctor B on female and male patients. Doctor B is less receptive to advice for male patients, so tr offers less advice albeit with higher confidence on the male population relative to the female population. This strategy leads tr to add more value compared to the other advisors to Doctor B, significantly improving decisions cost-effectively on the female population without significantly diminishing the value Doctor B produces when diagnosing male patients.
This case study represents a greater opportunity for adding value compared to the case study shown in the main results because Doctor B is receptive to advice on cases they would most benefit accepting advice for. Consequently, all advisors are able to add more value to this expert than they are able to add to Doctor A, however, tr still adds the most value to both experts relative to alternatives, demonstrating that tr is able to adapt its advising strategy to provide value-adding advice regardless of whether an expert’s ADB is calibrated to their strengths and weaknesses. When opportunities to improve on the expert’s behavior are limited, such as for the Doctor A, other methods cannot add value at all, while tr is able to add some value by offering only a few pieces of highly-confident and cost-effective advice. When there exist greater opportunities to improve on the expert’s behavior, tr adds the most value by offering more advice cost-effectively than it is able to when opportunities are limited.
C.5 Impact of discretion model accuracy on advising performance
ReV-AI advisors are optimized to account for human ADB by utilizing a model which estimates the human’s ADB. In all main results and case studies, we simulate real-world procedures by first learning a discretion model that estimates the human’s ADB. These models are always imperfect estimates of the humans’ true ADB. While imperfect, the models used for the main results were decent approximations of the human’s true ADB (AUC $>0.9$). In this section, we further analyze the value-added from tr’s advice as the human’s ADB deviates further from the discretion model used to optimize tr. It is critical to understand AI-advising performance under imperfect discretion estimates because, in practice, it may not be possible to learn an accurate discretion model. Further, it is critical to have mechanisms in place to ensure poor discretion models do not lead to advisors that can diminish human standalone value.
We conduct an experiment on the Heart Disease dataset in which we pair a trained tr advisor with different simulated humans. The first human exhibits the same ADB that was used to train tr, and thus the AUC of the discretion model evaluated on the human’s true ADB is $>0.9$. For the subsequent humans we add random noise to their ADB such that the discretion model used to train tr is less and less representative of the human’s true ADB. We include in the legend the AUC of the discretion model on each human’s ADB. Lower AUC values mean the human exhibits a true ADB that deviates more from what tr was trained on.
Results are shown in Figure 9. The solid lines show model performance after a robustness mechanism using validation data was applied, while the corresponding dashed lines of the same color demonstrate the result when the robustness mechanism is not applied. [^7] Note that in most cases, the dashed and solid line overlap, meaning the robustness mechanism did not activate to produce different results The robustness mechanism involves evaluating the advisor performance when paired with the human on a validation set of data, and, if the advisor is found to diminish the human’s value, the advisor is not used. In practice, assessing how the advisor performs with the individual expert on a set of low-stakes validation tasks is critical for producing reliable, value-adding advisors.
Intuitively, we find that as the human’s ADB deviates more from the model used to train tr, the value-added by tr diminishes. In some cases, we find that tr can begin to diminish the human’s value if opportunities for improvement are limited (such as at higher costs) and the robustness mechanism is not applied, such as in Figure 9b. With the robustness mechanism applied, the model would not be deployed if it diminishes the human’s standalone value, preventing possible losses, and demonstrating how critical it is to evaluate AI advising models on low-stakes validation data and ensure mechanisms are in place to prevent possible losses.
Figure 9: Results when an already trained tr advisor is paired with different humans that exhibit ground truth ADB that deviates from the ADB tr was trained on. AUC values in the legend represent AUC of tr’s estimated model of the human’s ADB on the human’s ground truth ADB, such that lower AUC values mean the human’s ADB deviates more from the model used to train tr. Solid lines show results of tr paired with the human while utilizing a robustness mechanism, while the dashed lines show the same pairing but without implementing the robustness mechanism. In many cases the lines overlap, meaning the mechanism does not have any effect. However, the robustness mechanism ensures that the advisor does not diminish the human’s value, as shown in the middle panel.